in the same above question  rectangular hexagon 
AB+AC+AD+AE+AF=6AO

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello mohit you havent attached a pic, but i ll still ans the ques :)
AB+AC+AD+AE+AF=AB+(AC+AF)+AD+AE
                                =AB+(AC+CD)+AD+AE   [since AF=CD]
                                =AB+AD+AD+AE
                                =2AD+(AB+AE)
                                =2AD+(ED+AE)               [since  AB=ED]
                                =2AD+AD
                                =3AD
                                =3*(2AO)          [since O is the center and AO=OD]
                                =6AO
kindly approve :)